A salesman earns Php 2500 a week plus 6% commission on weekly sales exceeding Php 75000. If his salary plus commission during one week amounted to Php 5500 how much was his sale?
Answer by TimothyLamb(4379)

y = 0.06(x - 75000) + 2500
5500 = 0.06(x - 75000) + 2500
5500 = 0.06x - 75000*0.06 + 2500
0.06x = 5500 + 75000*0.06 - 2500
x = (5500 + 75000*0.06 - 2500)/0.06
x = 125000
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---